OpenStudy (anonymous):

y=3x^-6x-5 ....trying yo find X intercept and can't.. please help.

OpenStudy (anonymous):

is it 3x^(-6x-5) or u have done the typing error

OpenStudy (anonymous):

oh!

OpenStudy (anonymous):

y=3x^2-6x-5

OpenStudy (anonymous):

let y=0 x=(3+2sqrt6) or x=(3-2sqrt6)/3 [by using quadratic]

OpenStudy (anonymous):

y=-5 (by leting x to b 0)

OpenStudy (anonymous):

the it means that (0;-5)and[(3-2sqrt6)/3;0]and[(3+2sqrt6)/3;0]

OpenStudy (anonymous):

ok. I got (0,-5) that is my y-intercept.....what would the x be?

OpenStudy (anonymous):

(-0.6;0) or(2.6)

OpenStudy (anonymous):

i've forgot 0 it should b (2.6;0)